Determine whether the following series converges or diverges. If the series converges, find its sum.

a. βˆ‘βˆžπ‘›=1 (1+2n+1)/3n

b. βˆ‘βˆžπ‘›=1 (3+2n)/(2n+2)

c. βˆ‘βˆžπ‘›=1 1/(4n2-1)

1
Expert's answer
2021-06-10T07:26:58-0400

a.


"\displaystyle\sum_{n=1}^{\infin}\dfrac{1+2^{n+1}}{3n}=\text{diverges}"

"\lim\limits_{n\to \infin}a_n=\lim\limits_{n\to \infin}\dfrac{1+2^{n+1}}{3n}=\infin\not=0"

The series diverges by the Test for divergence.


b.


"\displaystyle\sum_{n=1}^{\infin}\dfrac{3+2^{n}}{2^{n+2}}=\text{diverges}"

"\lim\limits_{n\to \infin}a_n=\lim\limits_{n\to \infin}\dfrac{3+2^{n}}{2^{n+2}}=\lim\limits_{n\to \infin}\dfrac{\dfrac{3}{2^{n}}+\dfrac{2^{n}}{2^{n}}}{\dfrac{2^{n+2}}{2^{n}}}"

"=\lim\limits_{n\to \infin}\dfrac{\dfrac{3}{2^{n}}+1}{4}=\dfrac{1}{4}\not=0"

The series diverges by the Test for divergence.


c.


"\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{4n^2-1}=\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{(2n-1)(2n+1)}"

"=\dfrac{1}{2}\bigg(\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{2n-1}-\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{2n+1}\bigg)""=\dfrac{1}{2}\bigg(\dfrac{1}{1}-\dfrac{1}{3}+\dfrac{1}{3}-\dfrac{1}{5}+...\bigg)"

"\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{4n^2-1}=\dfrac{1}{2}"



The series converges and its sum is "\dfrac{1}{2}."